# Vector valued estimates for matrix weighted maximal operators and product $\mathrm{BMO}$

### v2

@@ A.2 Convex-set valued functions in the complex setting
We denote by $\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$ the closed unit ball in $\mathbb{R}^{2d}$ and by $\mathbf{\overline{B}}_{\mathbb{C}^{d}}$ the closed unit ball in $\mathbb{C}^{d}$, so $\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=R(\mathbf{\overline{B}}_{\mathbb{C}^{d}})$.
A real ellipsoid (more precisely, a nondegenerate centrally symmetric ellipsoid) in $\mathbb{R}^{2d}$ is by definition a subset $E$ of $\mathbb{R}^{2d}$ of the form $E=A\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$ for some invertible real linear map $A:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$. In this case, the polar decomposition yields $E=(AA^{*})^{1/2}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$. It follows that for each ellipsoid $E$ in $\mathbb{R}^{2d}$ there is a unique real positive definite linear map $A:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ with $E=A\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$, which we denote by $A:=M_{\mathbb{R},2d}(E)$.
A complex ellipsoid (more precisely, a nondegenerate centrally symmetric ellipsoid) in $\mathbb{C}^{d}$ is by definition a subset $E$ of $\mathbb{C}^{d}$ of the form $E=A\mathbf{\overline{B}}_{\mathbb{C}^{d}}$ for some invertible complex linear map $A:\mathbb{C}^{d}\to\mathbb{C}^{d}$. In this case, the polar decomposition yields $E=(AA^{*})^{1/2}\mathbf{\overline{B}}_{\mathbb{C}^{d}}$. It follows that for each ellipsoid $E$ in $\mathbb{C}^{d}$ there is a unique complex positive definite linear map $A:\mathbb{C}^{d}\to\mathbb{C}^{d}$ with $E=A\mathbf{\overline{B}}_{\mathbb{C}^{d}}$, which we denote by $A:=M_{\mathbb{C},d}(E)$.
If $A:\mathbb{C}^{d}\to\mathbb{C}^{d}$ is any complex linear map, then under our identification of $\mathbb{C}^{d}$ with $\mathbb{R}^{2d}$, $A$ corresponds to the real linear map $RAR^{-1}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$. Concretely, $RAR^{-1}$ is obtained by $A$ through replacing each complex entry $a_{ij}$ of $A$ with
$H(a_{ij})=\begin{bmatrix}\mathrm{Re}(a_{ij})&-\mathrm{Im}(a_{ij})\\ \mathrm{Im}(a_{ij})&\mathrm{Re}(a_{ij})\end{bmatrix}.$
Observe that the map $H:\mathbb{C}\to\mathbb{R}^{2\times 2}$ is a ring monomorphism. We write $H(A):=RAR^{-1}$. It is then easy to see that if $A$ is complex hermitian, respectively unitary, respectively positive definite, then $H(A)$ is real symmetric, respectively orthogonal, respectively positive definite. In particular, if for each $z\in\mathbb{C}$, $L_{z}:\mathbb{C}^{d}\to\mathbb{C}^{d}$ denotes multiplication with $z$, then we can consider the corresponding linear map $\widetilde{L}_{z}:=RL_{z}R^{-1}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$. So for each $z\in\mathbb{C}$ and for each nonempty subset $E$ of $\mathbb{C}^{d}$ we have $zE=E$ if and only if $\widetilde{L}_{z}(R(E))=R(E)$. Moreover, if $z\in\mathbb{C}$ with $|z|=1$, then $\widetilde{L}_{z}$ is easily seen to be a real orthogonal map.
The precise relation between real and complex ellipsoids is explained in the following lemma.
@@
$\mathcal{C}_{d}(\mathbb{C})=\{R^{-1}(E):\leavevmode\nobreak\ E\in\mathcal{C}_{2d}(\mathbb{R}^{2d})\text{ and }\widetilde{L}_{z}(E)=E\}.$ (A.1)
@@
$E$ $=\widetilde{L}_{z}E=\widetilde{L}_{z}M_{\mathbb{R},2d}(E)\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=(\widetilde{L}_{z}M_{\mathbb{R},2d}(E)(\widetilde{L}_{z}M_{\mathbb{R},2d}(E))^{*})^{1/2}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$
$=\widetilde{L}_{z}M_{\mathbb{R},2d}(E)\widetilde{L}_{z}^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}},$
where we used the fact that $\widetilde{L}_{z}$ is an orthogonal linear map. Since $\widetilde{L}_{z}M_{\mathbb{R},2d}(E)\widetilde{L}_{z}^{-1}$ is positive definite, by the uniqueness of $M_{\mathbb{R},2d}(E)$ it follows that
$M_{\mathbb{R},2d}(E)=\widetilde{L}_{z}M_{\mathbb{R},2d}(E)\widetilde{L}_{z}^{-1},$
thus also
$M_{\mathbb{R},2d}(E)=\widetilde{L}_{z}^{-1}M_{\mathbb{R},2d}(E)\widetilde{L}_{z}.$
Now we set $\widetilde{E}:=R^{-1}(E)$ and compute
$\widetilde{E}=R^{-1}M_{\mathbb{R},2d}(E)\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=R^{-1}M_{\mathbb{R},2d}(E)R\mathbf{\overline{B}}_{\mathbb{C}^{d}}.$
In order to show that $\widetilde{E}\in\mathcal{C}_{d}(\mathbb{C})$, it suffices to show that the invertible real linear map $R^{-1}M_{\mathbb{R},2d}(E)R:\mathbb{C}^{d}\to\mathbb{C}^{d}$ is in fact complex linear. To this end, it suffices to prove that
$L_{z}R^{-1}M_{\mathbb{R},2d}(E)R=R^{-1}M_{\mathbb{R},2d}(E)RL_{z},$
for all $z\in\mathbb{C}$ in $|z|=1$. For such $z$ we compute
$L_{z}R^{-1}M_{\mathbb{R},2d}(E)RL_{z}^{-1}=L_{z}R^{-1}\widetilde{L}_{z}^{-1}M_{\mathbb{R},2d}(E)\widetilde{L}_{z}RL_{z}^{-1}$
$=L_{z}R^{-1}RL_{z}^{-1}R^{-1}M_{\mathbb{R},2d}(E)RL_{z}R^{-1}RL_{z}^{-1}$
$=R^{-1}M_{\mathbb{R},2d}(E)R.$
@@
We now show the inclusion $\subseteq$ in (A.1). Let $\widetilde{E}$ be any complex ellipsoid in $\mathbb{R}^{2d}$. We set $E:=R(\widetilde{E})$. Then, we compute
$E=RM_{\mathbb{C},d}(\widetilde{E})R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}.$
The map $RM_{\mathbb{C},d}(\widetilde{E})R^{-1}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}$ is invertible real linear, thus $E$ is a real ellipsoid in $\mathbb{R}^{2d}$. Moreover, for all $z\in\mathbb{C}$ with $|z|=1$ we have
$\widetilde{L}_{z}E$ $=\widetilde{L}_{z}RM_{\mathbb{C},d}(\widetilde{E})R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=RL_{z}R^{-1}RM_{\mathbb{C},d}(\widetilde{E})R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$
$=RL_{z}M_{\mathbb{C},d}(\widetilde{E})R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=RM_{\mathbb{C},d}(\widetilde{E})L_{z}R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=RM_{\mathbb{C},d}(\widetilde{E})L_{z}\mathbf{\overline{B}}_{\mathbb{C}^{d}}$
$=RM_{\mathbb{C},d}(\widetilde{E})\mathbf{\overline{B}}_{\mathbb{C}^{d}}=E,$
where we used the fact that $L_{z}M_{\mathbb{C},d}(\widetilde{E})=M_{\mathbb{C},d}(\widetilde{E})L_{z}$ due to the complex linearilty of $M_{\mathbb{C},d}(\widetilde{E})$ as well as the fact that the closed unit ball of $\mathbb{C}^{d}$ remains invariant under multiplication with $z$.
In particular, the complex ellipsoids in $\mathbb{C}^{d}$ form a closed subset of the real ellipsoids in $\mathbb{R}^{2d}$, for orthogonal maps induce isometries with respect to the Hausdorff metric. ∎
The previous exposition is related to [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.7]. The idea behind this result is the fact that for any measurable convex body valued function $F,$ the function $G(x)$ defined as the John ellipsoid of the convex body $F(x)$ is also measurable. For the reader’s convenience, we include the precise statement in the context of complex convex body valued functions.
Theorem 16 (Theorem 3.7 in [Cruz_Uribe_Bownik_Extrapolation]).
@@
The proof of this result is based on [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.8], [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.9] and [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.10], in a way analogous to that of the real convex body context. Since this is one of the crucial points to get Theorem A, we include the statements of these lemmata and their proofs.
Lemma 17 (Lemma 3.8 in [Cruz_Uribe_Bownik_Extrapolation]).
@@
Proof.
We follow the same approach as in [Cruz_Uribe_Bownik_Extrapolation]. For every $x\in\Omega,$ define $G(x)$ as the unique John ellipsoid of $F(x),$ which has nonempty interior. This already satisfies (A.2). It is only left to prove the measurability of $G.$
To this end, consider a dense sequence $P_{1},P_{2},\ldots$ of invertible matrices in $\mathrm{M}_{{d}}({\mathbb{C}}).$ In particular, we have that for any ellipsoid $E$ with nonempty interior, we can express $E=P\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}$ with $P\in\mathrm{M}_{{d}}({\mathbb{C}})$ and for any $\varepsilon>0$ there exists $n\geq 1$ such that
$P_{n}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}\subseteq E\subseteq(1+\varepsilon)P_{n}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}.$
We construct by induction a sequence of measurable convex body valued functions $G_{n}\colon\Omega\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ such that for every $n\geq 1,$ $G_{n}(x)$ is either an ellipsoid with nonempty interior or $G_{n}(x)=\{0\}$ for every $x\in\Omega.$ First, define
$G_{1}(x)=\begin{cases}P_{1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}&\text{ if }P_{1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}\subseteq F(x),\\ \{0\}&\text{ otherwise. }\end{cases}$
The values of $G_{1}$ are either an ellipsoid with nonempty interior or $\{0\}$ by construction, while it is easy to see that it is also measurable since the set
$\{x\in\Omega\colon P_{1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}\nsubseteq F(x)\}=\{x\in\Omega\colon F(x)\cap P_{1}(\mathbb{C}^{d}\setminus\mathbf{\overline{B}}_{\mathbb{C}^{{d}}})\neq\emptyset\}$
is measurable due to $F$ being measurable. Assume now that we have defined measurable functions $G_{1},\ldots,G_{n}$ and define
$G_{n+1}(x)=\begin{cases}P_{n+1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}&\text{if }P_{n+1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}\subseteq F(x)\text{ and }m_{d}(P_{n+1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}})>m_{d}(G_{n}(x)),\\ G_{n}(x)&\text{ otherwise.}\end{cases}$
Again, by construction $G_{n+1}(x)$ is either an ellipsoid with nonempty interior or just the trivial convex body $\{0\}.$ We need to see that $G_{n+1}$ is also measurable. To do so, remember that the volume functional $K\mapsto m_{d}(K)$ is a continuous mapping from $\mathcal{K}_{b}(\mathbb{C}^{d})$ to $[0,\infty),$ since this is the same as the volume functional $K\mapsto m_{2d}(K)$ from $\mathcal{K}_{b}(\mathbb{R}^{2d})$ to $[0,\infty).$ This and the measurability of $G_{n}$ implies that $m_{d}(G_{n}(x))\colon\Omega\rightarrow[0,\infty)$ is also a measurable mapping (see [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.5]). Now, for any given open set $U\subseteq\mathbb{C}^{d},$ we have that
$\{x$ $\in\Omega\colon G_{n+1}(x)\cap U\neq\emptyset\}$
=($=(\{x\in\Omega\colon m_{d}(G_{n}(x))\geq m_{d}(P_{n+1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}})\}$
$\cap\{x\in\Omega\colon G_{n}(x)\cap U\neq\emptyset\})$
($\cup(\{x\in\Omega\colon m_{d}(G_{n}(x))<m_{d}(P_{n+1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}})\}$
$\cap\{x\in\Omega\colon P_{n+1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}\subseteq F(x)\text{ and }P_{n+1}\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}\cap U\neq\emptyset\}).$
By the previous considerations, each of the sets appearing on the right-hand side is measurable, so $G_{n+1}$ is measurable as well.
The last step of the proof is to show that the sequence $G_{n}(x)$ converges to $G(x)$ in the Hausdorff metric. This will imply that $G$ is measurable with respect to the Hausdorff topology, which is equivalent to the definition that we have used of measurability of convex body valued functions (see [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.5]). Assume this is not the case for a given $x\in\Omega.$ By our construction, the density of the sequence $P_{1},P_{2},\ldots$ and the maximality of the John ellipsoid $G(x),$ we have that $m_{d}(G_{n})\to m_{d}(G(x))$ as $n\to\infty.$ The Blaschke selection theorem [Schneider_Convex_Bodie_1993, Theorem 1.8.7] asserts that any bounded sequence of convex bodies has a subsequence that converges to a convex body, so we have a subsequence $G_{n_{k}}(x)$ converging to an ellipsoid $E^{\prime}$ with $m_{d}(E^{\prime})=m_{d}(G(x))$ but with $E^{\prime}\neq G(x),$ which contradicts the fact that the John ellipsoid is unique. Note that $E^{\prime}$ is a complex ellipsoid, since real ellipsoids are closed in the Hausdorff topology and complex ellipsoids are a closed subset of the real ellipsoids. Thus $G_{n}(x)\to G(x)$ and the last function is also measurable. ∎
@@
Another approach consists in noting that the map sending each convex body with nonempty interior to its John ellipsoid is continuous with respect to the Hausdorff distance. In the real case, [DKPS2024, Subsection A.4.3] appeals directly to [Mordhorst_2017] for this result. An extension to the complex setting is performed in detail in [DKPS2024, Subsection A.4.7]. Combining this with the complex version of [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.5], we immediately deduce the complex version of [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.8], since the composition of measurable maps remains measurable.
The statement and proof of [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.9] for complex convex bodies correspond to Lemma 4 in the present work. The proof included in Section 2.7 already covers the case of complex convex bodies.
Finally, [Cruz_Uribe_Bownik_Extrapolation, Lemma 3.10] relates measurable complex ellipsoid valued mappings $G$ to measurable matrix valued mappings with mutually orthogonal columns.
Lemma 19 (Lemma 3.10 in [Cruz_Uribe_Bownik_Extrapolation]).
@@
The proof consists in constructing the columns $v_{1},\ldots,v_{d}$ of the matrix $W(x)$ as measurable mappings $\Omega\rightarrow\mathbb{C}^{d}$ and being mutually orthogonal (with some of them possibly null at a given $x\in\Omega$). Given our measurable mapping $G,$ take the measurable mapping $v_{1}\colon\Omega\rightarrow\mathbb{C}^{d}$ such that $|G(x)|=|v_{1}(x)|$ at every $x\in\Omega$ given by Lemma 4. Then define the mappings $J_{1}\colon\Omega\rightarrow\mathcal{K}(\mathbb{C}^{d})$ and $J_{1}^{\perp}\colon\Omega\rightarrow\mathcal{K}(\mathbb{C}^{d})$ given by $J_{1}(x)=\mathrm{span}\{v_{1}(x)\}$ and $J_{1}^{\perp}(x)=(J_{1}(x))^{\perp},$ respectively. By considering the field of Gaussian rationals $\mathbb{Q}(i)=\{p+iq\colon p,q\in\mathbb{Q}\},$ which is a dense set of $\mathbb{C},$ we get the sequence of selection functions $\{\lambda v_{1}(x)\}_{\lambda\in\mathbb{Q}(i)}$ for $J_{1},$ yielding immediately that the latter is measurable. A standard argument using the characterisations of measurability of closed set valued functions (see [Aubin_2009, Theorem 8.1.4]) shows that $J_{1}^{\perp}$ is also measurable.
Assume now that $v_{1},\ldots,v_{k}$ have already been defined for some $1\leq k<d.$ As before, define the mappings
$J_{k}(x)=\mathrm{span}\{v_{1}(x),\ldots,v_{k}(x)\},\qquad J_{k}^{\perp}(x)=(J_{k}(x))^{\perp},$
which are again measurable by the same standard arguments. Next, define the measurable mapping $G_{k}\colon\Omega\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d})$ taking values on the set of complex ellipsoids given by $G_{k}(x)=G(x)\cap J_{k}^{\perp}(x)$ (it might happen that $G_{k}(x)$ is the trivial ellipsoid). Applying Lemma 4, one gets a new measurable vector valued function $v_{k+1}$ such that it is orthogonal to $v_{1},\ldots,v_{k}.$
The vectors $v_{1}(x),\ldots,v_{d}(x)$ constructed in this way are the semiaxes of an ellipsoid, each one taken to be maximal in $G(x)$ at each step and given in decreasing order. Thus, these vectors are precisely the semiaxes of $G(x)$ and, if we define the measurable matrix valued function $W(x)$ whose columns are the previous vectors, we get the desired equality $G(x)=W(x)\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}.$
The proof of [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.7] follows by replacing $\mathbb{R}^{d}$ by $\mathbb{C}^{d}.$ Also, the real inner product and associated notions are replaced by the Hermitian product and respective notions. We include the details for completeness.
Proof of Theorem 3.7 in [Cruz_Uribe_Bownik_Extrapolation].
Note that if $F(x)$ is a convex body in $\mathbb{C}^{d}$ with nonempty interior for every $x\in\Omega,$ then Lemmata 17 and 19 yield the result. If that is not the case, one needs to divide $\Omega$ into sets $\Omega_{0},\ldots,\Omega_{d}$ such that $F(x)$ is contained in a linear subspace of dimension $k$ for every $x\in\Omega_{k}.$
Given the measurable mapping $F\colon\Omega\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{d}),$ define $J\colon\Omega\rightarrow\mathcal{K}(\mathbb{C}^{d})$ by
$J(x)=\mathrm{span}F(x)=\overline{\mathrm{conv}}\left(\bigcup_{\lambda\in\mathbb{Q}(i)}\lambda F(x)\right),$
which is measurable by a standard argument using the density of the Gaussian rationals in $\mathbb{C}.$ Also, the mapping $P\colon\Omega\rightarrow\mathrm{M}_{{d}}({\mathbb{C}})$ with $P(x)$ being the orthogonal projection from $\mathbb{C}^{d}$ to $J(x)$ at every $x\in\Omega$ is measurable as well (see [Cruz_Uribe_Bownik_Extrapolation, Theorem 3.6], which also holds in the complex setting by the previous discussion). Then, the sets
$\Omega_{k}=\{x\in\Omega\colon\mathrm{rank}(P(x))=k\},\qquad 0\leq k\leq d,$
are also measurable since the rank can be computed taking the determinants of minors. Now one only needs to construct the restrictions $W|_{\Omega_{k}}$ satisfying the conclusions of the theorem to get the desired function.
Let us fix $0\leq k\leq d.$ One can find measurable functions $v_{1},\ldots,v_{k}\colon\Omega_{k}\rightarrow\mathbb{C}^{d}$ such that $v_{1}(x),\ldots,v_{k}(x)$ are an orthonormal basis of $J(x)$ for every $x\in\Omega_{k}$ (see [Helson_Spectral_Theorem_1986, Theorem 2 in Section 1.3], which applies to complex Hilbert spaces). Denote the set of $s\times t$-matrices by $\mathrm{M}_{{s\times t}}({\mathbb{C}}).$ The matrix $M_{k}(x)\in\mathrm{M}_{{d\times k}}({\mathbb{C}})$ whose columns are $v_{1}(x),\ldots,v_{k}(x)$ is an isometry of $\mathbb{C}^{k}$ onto $J(x)$ for every $x\in\Omega_{k}$ with inverse given by $M_{k}^{\ast}(x)=(M_{k}(x))^{\ast}$ its conjugate transpose. Thus, we can consider the measurable function $F_{k}\colon\Omega_{k}\rightarrow\mathcal{K}_{\mathrm{bcs}}(\mathbb{C}^{k})$ given by $F_{k}(x)=M_{k}^{\ast}(x)F(x),$ which satisfies that $F_{k}(x)$ is a convex body with nonempty interior in $\mathbb{C}^{k}.$ Hence, Lemmata 17 and 19 yield a matrix valued function $W_{k}\colon\Omega_{k}\rightarrow\mathrm{M}_{{k}}({\mathbb{C}})$ such that its columns are mutually orthogonal and with
$W_{k}(x)\mathbf{\overline{B}}_{\mathbb{C}^{{k}}}\subseteq F_{k}(x)\subseteq\sqrt{d}W_{k}(x)\mathbf{\overline{B}}_{\mathbb{C}^{{k}}}$
for every $x\in\Omega_{k}.$
The actual mapping $W$ restricted to $\Omega_{k}$ is obtained as follows. Take $P_{k}$ the coordinate projection from $\mathbb{C}^{d}$ to $\mathbb{C}^{k},$ so $P_{k}\in\mathrm{M}_{{d}}({\mathbb{C}}).$ Then define $W(x)=M_{k}(x)W_{k}(x)P_{k}(x)$ for $x\in\Omega_{k}.$ This mapping is measurable and its columns are mutually orthogonal by construction. It is easy to check that it also satisfies the desired inclusions when applied to $\mathbf{\overline{B}}_{\mathbb{C}^{{d}}}.$
We now turn our attention to the second part of [Cruz_Uribe_Bownik_Extrapolation, Section 3], which concerns integrals of convex-set valued maps. The definitions of the Aumann integral and integrable bounded functions there are valid with $\mathbb{R}^{d}$ replaced by $\mathbb{R}^{2d}$ and thus also by $\mathbb{C}^{d}$.

### v3

@@ A.2 Convex-set valued functions in the complex setting
If $A:\mathbb{C}^{d}\to\mathbb{C}^{d}$ is any complex linear map, then $A$ corresponds to the real linear map $\widetilde{A}=RAR^{-1}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}.$ Now consider for each $z\in\mathbb{C}$ the map $L_{z}:\mathbb{C}^{d}\to\mathbb{C}^{d}$ denoting componentwise multiplication with $z.$ Then we can consider the corresponding linear map $\widetilde{L}_{z}\coloneq RL_{z}R^{-1}:\mathbb{R}^{2d}\to\mathbb{R}^{2d}.$ Observe that the matrix corresponding to $\widetilde{L}_{z}$ is a $(2d\times 2d)$-matrix with $d$ blocks of the form
$\begin{bmatrix}\mathrm{Re}(z)&-\mathrm{Im}(z)\\ \mathrm{Im}(z)&\mathrm{Re}(z)\end{bmatrix}$
on the diagonal. With the notation that we have introduced, we have that for each $z\in\mathbb{C}$ and for each nonempty $E\subseteq\mathbb{C}^{d}$ it holds that $zE=E$ if and only if $\widetilde{L}_{z}(R(E))=R(E).$ Moreover, if $z\in\mathbb{C}$ with $|z|=1,$ then $\widetilde{L}_{z}$ is a real orthogonal map. The next lemma states the precise relation between real and complex ellipsoids.
@@
$\mathcal{C}_{d}(\mathbb{C})=\{R^{-1}(E):\penalty 10000\ E\in\mathcal{C}_{2d}(\mathbb{R}^{2d})\text{ and }\widetilde{L}_{z}(E)=E\text{ for each }|z|=1\}.$ (A.1)
@@
Here, for a real ellipsoid $E$ in $\mathbb{R}^{2d}$ we denote by $M_{\mathbb{R}}(E)$ the unique positive semidefinite matrix with $E=M_{\mathbb{R}}(E)\mathbf{\overline{B}}_{\mathbb{R}^{2d}}.$ Likewise, for a complex ellipsoid $E$ in $\mathbb{C}^{d},$ we use $M_{\mathbb{C}}$ to denote the unique complex positive semidefinite matrix with $E=M_{\mathbb{C}}(E)\mathbf{\overline{B}}_{\mathbb{C}^{d}}.$
@@
$E$ $=\widetilde{L}_{z}E=\widetilde{L}_{z}M_{\mathbb{R}}(E)\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=(\widetilde{L}_{z}M_{\mathbb{R}}(E)(\widetilde{L}_{z}M_{\mathbb{R}}(E))^{*})^{1/2}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$
$=\widetilde{L}_{z}M_{\mathbb{R}}(E)\widetilde{L}_{z}^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}},$
because $\widetilde{L}_{z}$ is orthogonal, where in the third equality we have used the polar decomposition of $(\widetilde{L}_{z}M_{\mathbb{R}}(E))^{*}$ as in the proof of Lemma 14. Since $\widetilde{L}_{z}M_{\mathbb{R}}(E)\widetilde{L}_{z}^{-1}$ is positive semidefinite, uniqueness of $M_{\mathbb{R}}(E)$ implies that
$M_{\mathbb{R}}(E)=\widetilde{L}_{z}M_{\mathbb{R}}(E)\widetilde{L}_{z}^{-1}=\widetilde{L}_{z}^{-1}M_{\mathbb{R}}(E)\widetilde{L}_{z}.$
Now we need to check that $R^{-1}(E)$ is a complex ellipsoid. We compute
$R^{-1}E=R^{-1}M_{\mathbb{R}}(E)\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=R^{-1}M_{\mathbb{R}}(E)R\mathbf{\overline{B}}_{\mathbb{C}^{d}}.$
Thus, in order to show that $R^{-1}E\in\mathcal{C}_{d}(\mathbb{C}),$ it suffices to show that the real linear map $R^{-1}M_{\mathbb{R}}(E)R:\mathbb{C}^{d}\to\mathbb{C}^{d}$ is in fact complex linear. To this end, it suffices to prove that
$L_{z}R^{-1}M_{\mathbb{R}}(E)R=R^{-1}M_{\mathbb{R}}(E)RL_{z}$
for all $z\in\mathbb{C}$ with $|z|=1$. For such $z$ we compute
$L_{z}R^{-1}M_{\mathbb{R}}(E)RL_{z}^{-1}=L_{z}R^{-1}\widetilde{L}_{z}^{-1}M_{\mathbb{R}}(E)\widetilde{L}_{z}RL_{z}^{-1}$
$=L_{z}R^{-1}RL_{z}^{-1}R^{-1}M_{\mathbb{R}}(E)RL_{z}R^{-1}RL_{z}^{-1}$
$=R^{-1}M_{\mathbb{R}}(E)R.$
@@
We now show the inclusion $\subseteq$ in (A.1). Consider a complex ellipsoid $E$ in $\mathbb{C}^{d},$ so that $R(E)$ is a real ellipsoid in $\mathbb{R}^{2d}$ because
$R(E)=RM_{\mathbb{C}}(E)R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}.$
Moreover, for all $z\in\mathbb{C}$ with $|z|=1$ we have
$\widetilde{L}_{z}R(E)$ $=\widetilde{L}_{z}RM_{\mathbb{C}}(E)R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=RL_{z}R^{-1}RM_{\mathbb{C}}(E)R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}$
$=RL_{z}M_{\mathbb{C}}(E)R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=RM_{\mathbb{C}}(E)L_{z}R^{-1}\mathbf{\overline{B}}_{\mathbb{R}^{2d}}=RM_{\mathbb{C}}(E)L_{z}\mathbf{\overline{B}}_{\mathbb{C}^{d}}$
$=RM_{\mathbb{C}}(E)\mathbf{\overline{B}}_{\mathbb{C}^{d}}=R(E),$
where we used the fact that $L_{z}M_{\mathbb{C}}(E)=M_{\mathbb{C}}(E)L_{z}$ by linearity and the fact that the closed unit ball of $\mathbb{C}^{d}$ is invariant under multiplication with $z.$ This also shows that the complex ellipsoids in $\mathbb{C}^{d}$ form a closed subset of the real ellipsoids in $\mathbb{R}^{2d},$ since orthogonal maps induce isometries with respect to the Hausdorff metric. ∎
The previous exposition is related to [5, Theorem 3.7]. The idea behind this result is the fact that for any measurable convex body valued function $F,$ the function $G(x)$ defined as the John ellipsoid of the convex body $F(x)$ is also measurable. For the reader’s convenience, we include the precise statement in the context of complex convex body valued functions.
Theorem 16 (Theorem 3.7 in [5]).
@@
The proof of this result is based on [5, Lemma 3.8], [5, Lemma 3.9] and [5, Lemma 3.10], in a way analogous to that of the real convex body context. We include the statements of these lemmata without proofs, since they are the same as in the real context once one has Lemma 15. However, we mention some remarks about each of them.
Lemma 17 (Lemma 3.8 in [5]).
@@
As mentioned before, the proof in the complex case follows the same arguments as in the real case almost verbatim, only with the obvious modifications. It is worth mentioning that the proof makes use of the Blaschke Selection Theorem [39, Theorem 1.8.7], which is equally valid in the complex setting.
Also, this result can be shown by another approach. One can note that the map sending each convex body with nonempty interior to its John ellipsoid is continuous with respect to the Hausdorff distance. In the real case, [14, Subsection A.4.3] appeals directly to [33] for this result. An extension to the complex setting is performed in detail in [14, Subsection A.4.7]. Combining this with the complex version of [5, Theorem 3.5], we immediately deduce the complex version of [5, Lemma 3.8], since the composition of measurable maps remains measurable.
The statement and proof of [5, Lemma 3.9] for complex convex bodies correspond to Lemma 3 in the present work. The proof included in Section 2.4 already covers the case of complex convex bodies.
Finally, [5, Lemma 3.10] relates measurable complex ellipsoid valued mappings $G$ to measurable matrix valued mappings with mutually orthogonal columns.
Lemma 19 (Lemma 3.10 in [5]).
@@
The proof consists in constructing the columns $v_{1},\ldots,v_{d}$ of the matrix $W(x)$ as measurable mappings $\Omega\rightarrow\mathbb{C}^{d}$ and being mutually orthogonal (with some of them possibly null at a given $x\in\Omega$). This construction is the same for the complex case than for the real case. The major difference being only that, instead of using the rationals $\mathbb{Q}$ as a dense set of $\mathbb{R},$ one needs to take the field of Gaussian rationals $\mathbb{Q}(i)=\{p+iq\colon p,q\in\mathbb{Q}\}$ as a dense set of $\mathbb{C}.$
Once one has the complex setting versions of [5, Lemmata 3.8, 3.9 and 3.10], the proof of [5, Theorem 3.7] follows equally by replacing $\mathbb{R}^{d}$ by $\mathbb{C}^{d},$ and modifying the related notions accordingly (e.g. the Euclidean scalar product by the hermitian one).
We make some final remarks about the second part of [5, Section 3], which concerns integrals of convex-set valued maps. The definitions of the Aumann integral and integrable bounded functions there are valid with $\mathbb{R}^{d}$ replaced by $\mathbb{R}^{2d}$ and thus also by $\mathbb{C}^{d}$. Note that the definition of the Aumann integral makes use only of the topological space and real vector space structures of $\mathbb{R}^{d}$. Note also that complex symmetric sets are in particular real symmetric sets. These general principles can be applied through the rest of [5, Section 3] to obtain the complex versions of all results concerning the Aumann integral.